Factor the polynomial f(x). Then solve the equation f) o. 10) f(x) x3+5x2- 9x-45 State the domain of the rational function. (6 points) 11) g(x) =ー2 x +2 Given that the polynomial function has the given zero, find the other zeros. 12) f(x)=x3-4x2 + 9x-10:2
Answer #1

Solution : ( 10 )

f(x)=x^3+5x^2-9x-45

\mathrm{Factor}\:x^3+5x^2-9x-45:

\mathrm{Steps\::}

x^3+5x^2-9x-45

=\left(x^3+5x^2\right)+\left(-9x-45\right)

\mathrm{Factor\:out\:}x^2\mathrm{\:from\:}x^3+5x^2\mathrm{:\quad }x^2\left(x+5\right)

\mathrm{Factor\:out\:}-9\mathrm{\:from\:}-9x-45\mathrm{:\quad }-9\left(x+5\right)

=x^2\left(x+5\right)-9\left(x+5\right)

\mathrm{Factor\:out\:common\:term\:}\left(x+5\right)

=\left(x+5\right)\left(x^2-9\right)

=\left(x+5\right)\left(x^2-3^2\right)

\mathrm{Apply\:Difference\:of\:Two\:Squares\:Formula:\:}x^2-y^2=\left(x+y\right)\left(x-y\right)

=\left(x+5\right)\left(\left(x+3\right)\left(x-3\right)\right)

=\left(x+5\right)\left(x+3\right)\left(x-3\right)

\mathrm{We\:have\:now\:factored\:the\:polynomial\:completely.}

f(x)=\left(x+5\right)\left(x+3\right)\left(x-3\right)

\mathrm{Now\:that\:we\:have\:factored\:}f\left(x\right),\mathrm{\:the\:Zero\:Product\:Property\:tells\:us\:that\:the\:}

\mathrm{solutions\:to\:}f(x)=0\:\mathrm{\:are\:the\:values\:of\:}x\mathrm{\:that\:make\:a\:factor\:zero.\:So:}

x+5=0\mathrm{\:\:or\:\:}x+3=0\mathrm{\:\:or\:\:}x-3=0

\mathrm{Solving\:each\:of\:these\:for\:}x\mathrm{\:we\:get\::}

x=-5\mathrm{\:\:or\:\:}x=-3\mathrm{\:\:or\:\:}x=3

-------------------------------------------------------------------------------------------------

Solution : ( 11 )

g\left(x\right)=\frac{x-2}{x+2}

\mathrm{Domain\:of\:}\:\frac{x-2}{x+2}\::\quad

\mathrm{Steps\::}

\mathrm{Find\:undefined\:\left(singularity\right)\:points}:\quad

\mathrm{Take\:the\:denominator\left(s\right)\:of\:}\frac{x-2}{x+2}\mathrm{\:and\:compare\:to\:zero}

\mathrm{Solve\:}\:x+2=0:\quad

x+2=0

\mathrm{Subtract\:}2\mathrm{\:from\:both\:sides}

x+2-2=0-2

\mathrm{Simplify}

x=-2

\mathrm{The\:following\:points\:are\:undefined}

x=-2

\mathrm{The\:function\:domain}

x<-2\quad \mathrm{or}\quad \:x>-2

\mathrm{Hence,}

\mathrm{Domain\:of\:}\:\frac{x-2}{x+2}\::\quad \begin{bmatrix}\mathrm{Solution:}\:&\:x<-2\quad \mathrm{or}\quad \:x>-2\:\\ \:\mathrm{Interval\:Notation:}&\:\left(-\infty \:,\:-2\right)\cup \left(-2,\:\infty \:\right)\end{bmatrix}

-------------------------------------------------------------------------------------------------

Solution : ( 12 )

f(x)=x^3-4x^2+9x-10

\mathrm{Steps \::}

x^3-4x^2+9x-10=0

\mathrm{Solve\:by\:factoring}

\mathrm{Factor\:}x^3-4x^2+9x-10:

x^3-4x^2+9x-10

\mathrm{Use\:the\:rational\:root\:theorem}

a_0=10,\:\quad a_n=1

\mathrm{The\:dividers\:of\:}a_0:\quad 1,\:2,\:5,\:10,\:\quad \mathrm{The\:dividers\:of\:}a_n:\quad 1

\mathrm{Therefore,\:check\:the\:following\:rational\:numbers:\quad }\pm \frac{1,\:2,\:5,\:10}{1}

\frac{2}{1}\mathrm{\:is\:a\:root\:of\:the\:expression,\:so\:factor\:out\:}x-2

\mathrm{Compute\:}\frac{x^3-4x^2+9x-10}{x-2}\mathrm{\:to\:get\:the\:rest\:of\:the\:eqution:\quad }x^2-2x+5

(z-2) (2.2-22, +5)

\left(x-2\right)\left(x^2-2x+5\right)=0\quad------\rightarrow (1)

\mathrm{Using\:the\:Zero\:Factor\:Principle:}

From (1);

\mathrm{Solve\:}\:x-2=0:\quad x=2

\mathrm{and}

\mathrm{Similarly.\:\:From\:\:(1)};

\mathrm{Solve\:}\:x^2-2x+5=0:\quad

x^2-2x+5=0

Solve with the quadratic formula

For a quadratic equation of the form ax2 + bx + c = 0 the solutions are

x_{1,\:2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a}

\mathrm{For\:}\quad a=1,\:b=-2,\:c=5:\quad x_{1,\:2}=\frac{-\left(-2\right)\pm \sqrt{\left(-2\right)^2-4*1*5}}{2*1}

x_{1}=\frac{-\left(-2\right)+ \sqrt{\left(-2\right)^2-4*1*5}}{2*1}

=\frac{2+\sqrt{16}i}{2}

=\frac{2+4i}{2}

=\frac{2\left(1+2i\right)}{2}

=1+2i

\mathrm{and}

x_{2}=\frac{-\left(-2\right)- \sqrt{\left(-2\right)^2-4*1*5}}{2*1}

=\frac{2-\sqrt{16}i}{2}

=\frac{2-4i}{2}

=\frac{2\left(1-2i\right)}{2}

=1-2i

The final solutions to the quadratic equation are

x=1+2i,\:x=1-2i

The final solutions to the equation are:

x=2,\:x=1+2i,\:x=1-2i

\mathrm{\:Hence,\:\:the\:other\:two\:zeros}

x=1+2i,\:x=1-2i